Shift Summary for Eve shift, 29/Jul/99

Miguel Barrio (mbarrio@hep.uchicago.edu)
Fri, 30 Jul 1999 00:07:19 -0500

Shift Summary for Eve shift, 29/Jul/99

On Shift: Jim Graham, Miguel Barrio

Hours of beam available at all intensities: 7.5
Hours of beam available at nominal intensity: 7.5
Hours of E832 data (special runs not included): 7.0
Hours of special runs (laser scans, muon runs...): 0.0

Accounting of lost time:
Had to reboot atev2 and restart all SDAQ processes on all atev machines

========================================================================
Runs ended this shift:
Run 14060 - 15.3M events - BEAM_832

Tapes brought to FCC this shift: none

========================================================================
Accesses during shift: none

========================================================================
Overview of shift:
E832 datataking

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------

14060 15:30 Previous shift hands shift duties to us.
We are getting beam and everything is.

16:00 The onsplit daemons got messed up (probably due
to the reboot of all ktev machines). Basically,
what happened was that a tape drive was already
allocated when the daemon tried to allocate it
itself, thus resulting in a conflict. Sidney
logged on remotely and fixed the problem.

16:40 Beam's gone, MI vacuum problems.

16:40 Beam's back.

17:22 MCR calls (Steven). NMU2 tripped due to
overheating. Beam's gone.

17:30 Beam's back, magnet cooled down real fast
(Brisk, baby!)

18:35 Online kumacs are dandy...

18:50 Beam's gone, no message indicating why.

19:20 Dave N. changes He tubes to #5 and #6.

19:23 Beam's back.

20:40 SDAQ alarm: Etot_Temp LV Chan 1 reads .694V
whereas nominal is .8V. Paged J Hamm, he says
we should keep an eye on it for now.

20:42 Etot_Temp button goes green. But goes back to
red a couple of minutes later.

20:50 atev2's mouse is not responding. Check all
connections, the problem is at the machine.
Called Theo and Ashkan, wwho said we should
reboot.

21:19 Gas checklist is done.

14060 358 21:25 End 14060 to reboot atev2 and restart all SDAQ
processes. Did that and everything came back up
with all mice working.

14061 21:35 Start of run 14061. Etot_Temp problem persists
but hasn't gotten worse, the values are .695V
to .685V, acillating between these two.

========================================================================
Follow up Notes on Previous Shift

========================================================================
Items to Follow up on Next Shift

Keep an eye on Etot_Temp on the SDAQ window. If it goes lower than .65V,
page Jason Hamm.

Do not take a muon run... repeat with me "I will not take a muon run."
Good. (this message has been brought to you as per instructions from the
Run Tzar Mista R S Kessler).